# Questions: A Taxonomy for Critical Reflection in Machine-Supported Decision-Making

### v1

@@ 2.4. Reflection and Responsible Decision-Making
Reflection proved to be the most effective form of cognitive intervention for increasing cognitive engagement, among other methods such as checklists or feedback (Lambe et al., 2016; Prakash et al., 2019). Reflection, also referred to as reflective practice (Schön, 1983), or reflective thought, is a multi-faceted concept with multiple meanings (Nguyen et al., 2014; Rogers, 2001; Mann et al., 2009). We adopt the definition of the American philosopher John Dewey (1933), who defines reflective thought as the “active, persistent and careful consideration of any belief or supposed form of knowledge in light of the grounds that support it and further conclusion to which it tends” (p9). Consequently, reflection can promote reasoning, critical thinking and problem-solving, and thus improve decision-making (Schmidt and Mamede, 2023; Mamede and Schmidt, 2023; Lambe et al., 2016; Prakash et al., 2019; Walger et al., 2016; Ghanizadeh, 2017; Khoshgoftar and Barkhordari-Sharifabad, 2023; Hess et al., 2015). While reflection can occur before, during or after an action (Rogers, 2001), we focus on reflection during decision-making, i.e., reflection-in-action (Schön, 1983).
In areas with far-reaching decisions, such as law or healthcare, decision-makers have a professional responsibility, such as physicians who are committed to the well-being of patients. This professional responsibility involves an epistemological responsibility, which means that decision-makers are responsible for gathering and evaluating relevant information, and knowing the reasons for a particular course of action (van Baalen and Boon, 2015). In regard to past actions, knowing and providing reasons for an action allows for backward-looking responsibility or accountability. More importantly, however, in regard to future actions, knowing the reasons allows to take forward-looking responsibility by determining and directing actions to achieve a desired state (Loughran, 1996). By making it possible to be aware of the reasons for a decision, to formulate and weigh them up accordingly, and to change them if necessary (Mezirow, 1990), reflection in individual cases, i.e., reflection-in-action, can increase epistemological responsibility, both in a backward-looking and, more importantly, in a forward-looking sense. Ideally, this forward-looking responsibility leads to fairer and better outcomes for all parties involved (Ueda et al., 2024). Reflection-in-action could, for example, prevent a physician from jumping to conclusions after seeing seemingly similar patient cases throughout the day, which could lead to tunnel vision (Mamede and Schmidt, 2023). Consequently, to cite Dewey again, reflection “converts action that is merely appetitive, blind and impulsive into intelligent action” (Dewey, 1933, p17).
@@ 2.5. The Socratic Questioning Method
A common technique for facilitating reflection is to ask questions (Osmond and Darlington, 2005). One questioning technique that is widely used in education is the Socratic method. The Socratic method consists of systematically asking questions that help to clarify concepts, improve understanding, and uncover gaps in knowledge (Paul, 2019). Further, it helps to discover own thoughts, to analyse assumptions, information and inferences, and to arrive at own judgements through own reasoning (Paul, 2019). Besides, answering questions promotes reasoning (Alfano, 2019). Socratic questions are thus closely linked to critical thinking, which aims at “judging in a reflective way what to do or what to believe” (Facione, 2000).
@@
In general, independent thinking can promote the decision-autonomy of the decision-maker, which in turn can affect their motivation and well-being (Ryan and Deci, 2000). Findings from educational research suggest that motivation increases cognitive engagement (Rotgans and Schmidt, 2011; Blumenfeld et al., 2006) and even functions as a antecedent to it (Singh et al., 2022). As soon as students have the opportunity to participate in the organisation of learning activities, their interest and motivation increase (Blumenfeld et al., 2006). Without claiming a linear causality, questions seem to encourage reflection, which can increase decision-making autonomy, which can increase motivation, which can increase engagement.
In view of this, and the connection between critical reflection and reasoning and the associated positive effect on decision-making, we see potential in the transfer of questions, in particular through the Socratic method, from the education sector to the context of machine-supported decision-making. The Socratic method is slowly gaining traction in the field of human-machine collaboration, with a conceptual proposal of a virtual assistant that promotes reflection through Socratic questions (Lara and Deckers, 2020), and an educational chatbot based on a large language model and the Socratic method, which was found to improve critical reflection compared to traditional chatbots (Favero et al., 2024). There is, however, no systematic approach as to how the Socratic method and questions for critical reflection in general can be transferred to the area of machine-supported decision-making. This is partially because the Socratic method is primarily intended for the debate of philosophical issues. As such, a current taxonomy for Socratic questions (Paul, 2019, 2010) is primarily aimed at teachers and students to probe and foster thinking. We therefore aim to address this gap by transferring this taxonomy for Socratic questions and combining it with other work. In doing so, we concretise aforementioned conceptual proposals for human-machine collaborations that go beyond the traditional paradigm of machine recommendations and explanations (Ma et al., 2024b; Zhang et al., 2023; Haselager et al., 2023; Schmidt and Mamede, 2023).
@@ 3. Methodology
In the following, we will discuss how we have synthesised three earlier works in order to derive our question taxonomy, namely 1) a taxonomy of Socratic questions (Paul, 2019), 2) a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022), and 3) a revision of Bloom’s taxonomy for categorising educational objectives (Krathwohl, 2002; Forehand, 2010).
In the first step, we determine the elements for reflection, i.e., what the question shall relate to. To do so, we derive the overall structure of our taxonomy from a taxonomy for Socratic questions (Paul, 2019), which divides questions into the following components of reasoning:
  • Purpose: Questions that relate to the goal or objective.
  • Question at issue: Questions that relate to the problem or issue that gave rise to the question.
  • Information: Questions that relate to background information, such as data, facts, observations or experiences.
  • Interpretation and Inferences: Questions that relate to how meaning was derived and conclusions drawn.
  • Concepts: Questions that relate to underlying theories, definitions, or models that define a thought.
  • Assumptions: Questions that relate to taken for granted presuppositions of a thought.
  • Implications and Consequences: Questions that focus on implications.
  • Point of View: Questions that relate to the frame of reference or perspective of a thought.
We translate these categories and apply them to the context of decision-making in which operators use DSS. As such we the derive the following question categories: case information, e.g., input data, (information), relevance of data (information, interpretation and inference), dataset (concepts), causal structure of recommendation (interpretation and inferences), alternatives to recommendation (question, purpose), assumptions and expectations of decision-maker (assumptions, interpretation and inferences), stakeholder preferences (point of view), consequences of recommendation (implications), what-to-be that outcome (purpose), and model behaviour (assumptions) (see column ‘Element for Reflection’ in table 1).
In the second step, we identify elements, such as data or model behaviour through explanations, that can serve as information for the basis of the content of questions, i.e., how the questions can be enriched. To make appropriate use of the information derived from XAI, and given the various techniques available, we turn to a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022). The authors cluster more than 50 common questions operators have towards a DSS into the following categories:
@@
  • How to be That (a different prediction): Questions about how to change
@@
The authors use these question categories to map them to existing XAI techniques, such as feature importance or counterfactuals. This mapping is intended to help designers and developers select the appropriate XAI technique to provide the relevant answer, i.e., explanation.
In view of the overlap between the categorisation of Socratic questions and the common questions operators have, we use the mapping of questions to XAI techniques provided by the authors of (Liao et al., 2020, 2021; Liao and Varshney, 2022). Similarly, we assign XAI techniques to our identified question types in order to enrich possible questions with information (column ‘Useful Information’ in table 1). On the one hand, this means that (some) explanations can be (re-)formulated as questions (Danry et al., 2023), e.g., “Does outcome $Y$ follow from feature $x$?”. On the other hand, questions can relate to the decisive information extracted from explanations, e.g., “Is feature $x$ the relevant factor to focus on?”. So compared to the XAI question bank (Liao et al., 2020, 2021; Liao and Varshney, 2022), we reverse the process by repurposing explanations in order to ask the operators questions that stimulate critical reflection, instead of starting with questions from operators in order to arrive at possible explanations.
In the final step, and since questions can stimulate different thinking processes, we draw inspiration, albeit only marginally, from a revised version of Bloom’s taxonomy in order to capture the scope of questions (Krathwohl, 2002; Forehand, 2010; Anderson and Krathwohl, 2001). In Bloom’s taxonomy, questions are categorised hierarchically into six levels of cognitive processes, which are: remembering, understanding, applying, analysing, evaluating, and creating. The last three levels, analysing, evaluating, and creating, are also components of critical thinking (Paul, 2019). We thus hypothesise that in order to effectively stimulate critical reflection, questions should address these three higher-level cognitive abilities, where:
@@
This distinction is meant to help formulate concrete questions for specific use cases.
@@ 4. Taxonomy of Questions for Critical Reflection
In order to help create fruitful questions for critical reflection, we present our taxonomy of question types (table 1). We refer to table 2 for some example questions from the medical domain, where a physician uses a DSS in order to find a diagnosis or treatment option for a patient. For this, a collection of general clinical questions served as input (Ely, 2000).
Table 1. Our question taxonomy, which supports the systematic creation of questions that stimulate critical reflection during decision-making. From a taxonomy of Socratic questions we take the elements of thought (Paul, 2019), indicated in italic, and translate them to the decision-making process between human and DSS. We provide helpful information that can enrich the questions. To map existing XAI techniques to question types, we utilise a question bank for human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022). IDs marked with an asterisk (*) indicate questions that address the level of creating, according to Bloom’s categorisation of questions for education (Krathwohl, 2002). We refer to table 2 for sample questions.
ID Element for Reflection (What?) Description (Why?) Useful Information (How?)
@@
Questions to help further assess, inspect, scrutinise, and contextualise data points, to ensure its quality and reliability. Input Data (e.g., tabular patient data)
@@
(Information, Interpretation and Inference)
Questions to help evaluate data to ensure that it is valid by adequately supporting the hypothesis / recommendation. Derived explanations (XAI) can be framed as questions. Feature Contribution (e.g., SHAP (Lundberg and Lee, 2017), LIME (Ribeiro et al., 2016))
@@
(Concepts )
Questions to help inspect, assess and judge the assumptions built into the model to ensure that data adequately represents the phenomenon and that the conclusions derived are valid. Training Data, Datasheets (Gebru et al., 2021), Model Cards (Mitchell et al., 2019), FactSheets (Arnold et al., 2019), Global Feature Importance (Lundberg et al., 2020)
@@
(Interpretation and Inference)
Questions that help evaluate whether the outcome follows from data to ensure that the causal structure of the model / recommendation is sound. Can entail the previous questions of checking whether the data (Q1&Q2) and concept (Q3), i.e., model, are reliable and valid. Feature Contribution (e.g., (Lundberg and Lee, 2017), (Ribeiro et al., 2016))
@@
Questions that help to consider other possibilities and create other options for actions, to ensure that (larger) solution space is considered. Contextual Information, Differential diagnosis, Patient Preferences (Q7)
@@
Questions that elicit taken for granted assumptions to ensure that the decision-maker is aware of their reasons and can direct their actions accordingly. General Questions
@@
Questions that help the decision-maker enquire and take into account the preferences or needs of the people concerned (e.g., patient), in order to ensure a better or broader understanding of the problem at hand. General questions
Q8* Consequences of Recommendation
Questions that help elicit anticipation and forward-looking responsibility to ensure that (unintended) consequences are considered and mitigated. General questions
Q9* What to be that Outcome
Questions to support the exploration of alternative courses of action, by gaining insights from hypothetical scenarios derived from the DSS. Feature Perturbation, Counterfactuals (Karimi et al., 2021; Poyiadzi et al., 2020)

### v2

@@ 2.4. Reflection and Responsible Decision-Making
Amongst various forms of increasing cognitive engagement, such as checklists or feedback, stimulating reflection proved to be the most effective (Lambe et al., 2016; Prakash et al., 2019). While reflection can occur before, during or after an action (Rogers, 2001), we focus on reflection during decision-making, i.e., reflection-in-action (Schön, 1983). As mentioned earlier, reflection can promote reasoning, critical thinking and problem-solving, and thus improve decision-making.
In areas where decisions have far-reaching consequences, such as law or healthcare, decision-makers have a professional responsibility; physicians, for example, are committed to the well-being of patients. This professional responsibility involves an epistemological responsibility, which means that decision-makers are responsible for gathering and evaluating relevant information, and knowing the reasons for a particular course of action (van Baalen and Boon, 2015). In regard to past actions, knowing and being able to provide reasons for an action allows for backward-looking responsibility or accountability. More importantly, however, knowing the reasons makes it possible to determine and direct future actions in order to achieve a desired state, and thus to take forward-looking responsibility (Loughran, 1996). Reflection-in-action can therefore increase professional responsibility, both in a backward-looking and, more importantly, in a forward-looking sense. Ideally, this forward-looking responsibility leads to fairer and better decision outcomes for all parties involved (Ueda et al., 2024). Reflection-in-action could, for example, prevent a physician from jumping to conclusions after seeing seemingly similar patient cases throughout the day, which could lead to tunnel vision (Mamede and Schmidt, 2023). Reflection makes it possible to distinguish between relevant and irrelevant information, and, to quote Dewey again, “converts action that is merely appetitive, blind and impulsive into intelligent action” (Dewey, 1933, p.17).
@@ 2.5. The Socratic Questioning Method
A common technique to facilitate reflection is to ask questions (Osmond and Darlington, 2005). Moreover, questions encourage independent thinking, which can increase confidence in one’s own decision-making, i.e., decision self-efficacy, which in turn can affect the motivation and well-being of the decision-maker (Ryan and Deci, 2000). Findings from educational research suggest that motivation functions as an antecedent to cognitive engagement and also increases it (Rotgans and Schmidt, 2011; Blumenfeld et al., 2006; Singh et al., 2022). Questions thus seem to promote reflection, which can increase decision self-efficacy and motivation, which in turn can increase cognitive engagement.
To formulate relevant and effective questions, we want to point out two works, which we also use for our purposes. First, Bloom’s taxonomy for educational objectives (Krathwohl, 2002; Anderson and Krathwohl, 2001) categorises cognitive processes into six levels, which are: remembering, understanding, applying, analysing, evaluating, and creating. This categorisation is meant to help educators to define learning goals, exercises, and questions accordingly. To stimulate reflection, the last three levels, analysing, evaluating, and creating, seem the most promising as they are also components of critical thinking (Paul, 2019), which aims at “judging in a reflective way what to do or what to believe” (Facione, 2000). Second, the Socratic method is a questioning method that is widely used in education to develop and promote critical thinking skills (Makhene, 2019; Ho et al., 2023; Dalim, 2022; Torabizadeh et al., 2018). The Socratic method consists of systematically asking questions that help to clarify concepts, improve understanding, and uncover gaps in knowledge (Paul, 2019). Further, it helps to discover own thoughts, to analyse and evaluate assumptions, information and inferences, and to arrive at judgements through own reasoning (Paul, 2019). To help teachers to explore and encourage their students’ thinking through guided questions, a taxonomy of Socratic questions has been developed, defining eight types of questions that probe different parts of thinking, such as assumptions, reasons, or implications (Paul, 2019).
The Socratic questioning method is slowly gaining traction in the field of human-machine interaction. Based on the taxonomy of Socratic questions, a dataset of 110k questions that could be used in counselling and coaching is proposed, including a generation method of prompt-tuning a language model (Ang et al., 2023). The resulting questions fall into one of the six categories defined by Paul (2019): 1) clarification (“What do you mean by…?”), 2) probing assumptions (“Why do you assume…?”), 3) probing reasons (“How do you know that…?”), 4) probing implications (“If…, what is likely to happen?”), 5) probing alternatives (“What else should we consider?”), and 6) others that do not conform to the Socratic categories (Ang et al., 2023). In addition to this work, a concept of a virtual assistant that fosters moral enhancement through Socratic questions is proposed (Lara and Deckers, 2020), as well as chatbots based on language models that utilise the Socratic method for thought-provoking tutoring (Favero et al., 2024; Liu et al., 2024).
While these works appear promising, there is currently no systematic approach to using questions to promote critical reflection in decision-making, especially in combination with decision-support systems. Given the link between critical reflection and reasoning and the associated positive effect on decision-making, we see potential for the use of questions to promote decision self-efficacy and cognitive engagement. We therefore aim to address this gap by transferring an existing taxonomy of Socratic questions (Paul, 2019, 2010) to the context of machine-assisted decision-making. In doing so, we also aim for a higher level of granularity as compared to Ang et al. (2023), who utilise parts of the Socratic questions taxonomy in unchanged form to generate counselling and coaching questions. With our proposed taxonomy we provide a systematic way to formulate relevant questions for critical reflection during decision-making, based on concepts of critical thinking and the Socratic questioning method.
@@ 3. Methodology
In the following, we will discuss how we constructed our question taxonomy, which we did in two steps: identifying possible questions, and identifying information for formulating questions. Overall, we synthesised three prior works, namely 1) a taxonomy of Socratic questions (Paul, 2019), 2) a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022), and 3) a revision of Bloom’s taxonomy for educational objectives (Krathwohl, 2002; Forehand, 2010). The creation of the taxonomy resembles affinity diagramming with a deductive (top-down) approach, for which the Socratic questions functioned as a starting point and the XAI question bank helped us to translate the categories to the context of machine-assisted decision-making (Figure 1).
3.1. Phase 1: What to question?
In the first step, we identified relevant aspects of machine-assisted decision-making to which potential questions should relate, i.e.,what elements to reflect on. For this purpose, we used a taxonomy for Socratic questions based on critical thinking concepts (Paul, 2019, 2010). As Elder and Paul (2010) note:
“Whenever we think we think for a purpose within a point of view based on assumptions leading to implications and consequences. We use data, facts and experiences to make inferences and judgements based on concepts and theories to answer a questions or solve a problem.”
Accordingly, the authors categorise questions as follows into elements of thought or reasoning that allow to question the structure of thinking (Paul, 2019, 2010):
  • Purpose: Questions about the agenda, goal or objective.
  • Question at Issue: Questions about the problem or issue that gave rise to the question.
  • Information: Questions about background information, such as data, facts, observations or experiences.
  • Interpretation and Inferences: Questions about how meaning was derived and conclusions drawn.
  • Concepts: Questions about underlying theories, definitions, or models that define a thought.
  • Assumptions: Questions about taken for granted presuppositions of a thought.
  • Implications and Consequences: Questions about implications.
  • Point of View: Questions about the frame of reference or perspective of a thought.
While these elements of thought are helpful in general, the provided example questions (Table A1 in the Appendix) cannot be directly applied to the domain of decision-making (with DSSs). This is because the taxonomy of Socratic Questions is primarily intended for teachers and students in the classroom. We thus took the elements of thought as a starting point for our taxonomy and mapped them to the context of machine-assisted decision-making. The element Information, for example, can be mapped, as is also mentioned in the definition, to the dataset the DSS is trained on, as well as data of a specific (patient) case. To further specify relevant elements or aspects of machine-assisted decision-making, we used the XAI question bank, in which the authors cluster more than 50 common questions operators have towards a DSS into the following categories (Liao et al., 2020, 2021; Liao and Varshney, 2022):
@@ 3.1. Phase 1: What to question?
  • How to be That (a different prediction): Questions about how to change an instance to get a different outcome.
@@
Questions from the XAI question bank are about the logic of the model, its performance, the reasons for a prediction, possible changes, the data, and the output.
This leaves us with, on the one hand, the categorisation of questions based on the structure of thought, and on the other hand, the components of machine-assisted decision-making represented by clusters of common questions in the context of explainable AI (Figure 1). Through affinity diagramming we arrived at the following 10 elements: case information, relevance of data, dataset, causal structure of recommendation, alternatives to recommendation, assumptions of decision-maker, stakeholder preferences, consequences of decision, change intervention, and model behaviour.
Figure 1. On the left are the elements of thought based on the taxonomy for Socratic questions (Paul, 2019). On the right are the question categories identified in the XAI question bank (Liao et al., 2020). Based on these, we define our elements for reflection in machine-assisted decision-making in the middle column.
To define our final question categories, we followed an iterative process. We presented 10 initial categories, including some sample questions, to one orthopaedic surgeon and one senior researcher in orthopaedics who confirmed the completeness of the derived question types. We then refined the categories, without adding or removing any, expanded details, added descriptions, and added sample questions to a separate table (Table A3).
3.2. Phase 2: How to enrich questions?
After identifying what to question, the next step was to define how to derive and enrich specific questions of said categories. For this we again used the XAI question bank. For our purposes, however, we reversed the process. Instead of starting with questions from operators in order to arrive at suitable explanations, we use XAI techniques to extract information which then inform possible questions (column ‘Useful Information’ in Table 1). On the one hand, this means that (some) explanations can be (re-)formulated as questions (Danry et al., 2023), e.g., “Does outcome $Y$ follow from information $x$?”. On the other hand, questions can relate to information extracted from explanations, e.g., “Is information $x$ the relevant factor to focus on?”.
Finally, to capture the scope of questions, we used, albeit only marginally, a revised version of Bloom’s taxonomy (Krathwohl, 2002; Forehand, 2010; Anderson and Krathwohl, 2001). In Bloom’s taxonomy, questions are categorised hierarchically into six levels of cognitive processes, which are: remembering, understanding, applying, analysing, evaluating, and creating. The last three levels, analysing, evaluating, and creating, are also components of critical thinking (Paul, 2019). In order to effectively stimulate critical reflection, we thus hypothesise that questions should address these three higher-level cognitive abilities, where:
@@ 3.2. Phase 2: How to enrich questions?
This distinction, with the associated verbs, is meant to help formulate concrete questions for specific use cases.
@@ 4. Taxonomy of Questions for Critical Reflection
In the following, we will present our taxonomy, which consists of 10 question types or elements for reflection (Table 1). In order to demonstrate how to apply our taxonomy, we provide some sample questions from the medical and educational domain in Table A3 and Table A4 respectively.
Table 1. Our proposed taxonomy that supports the systematic creation of questions to stimulate critical reflection in machine-assisted decision-making. From a taxonomy of Socratic questions we take the elements of thought (Paul, 2019), indicated in italic, and map them to the context of machine-assisted decision-making. To do so and to enrich questions, we utilise and repurpose existing XAI techniques by reversing the mapping between questions and explanations in a question bank for human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022). IDs marked with an asterisk (*) indicate questions that address the level of creating, according to Bloom’s categorisation of questions for education (Krathwohl, 2002). For sample questions, we refer to Table A3 and Table A4 in the Appendix.
ID Element for Reflection (What?) Description Useful Information (How?)
@@
Questions that help further assess, inspect, and contextualise data points to ensure its quality, reliability, and completeness. Input Data (e.g., tabular patient data), Datasheets (Gebru et al., 2021)
@@
(Information, Interpretation & Inference)
Questions that help evaluate whether the data adequately supports the hypothesis / recommendation. Feature Contribution (e.g., SHAP (Lundberg and Lee, 2017), LIME (Ribeiro et al., 2016))
@@
Questions that help analyse and assess the data and to ensure it adequately represents the phenomenon (by highlighting limitations or characteristics of the dataset). Training Data, Datasheets (Gebru et al., 2021), FactSheets (Arnold et al., 2019)
@@
(Interpretation & Inference)
Questions that help evaluate whether the outcome follows from data to ensure that the causal structure of the model / recommendation is sound. Feature Contribution (e.g., (Lundberg and Lee, 2017), (Ribeiro et al., 2016)), Counterfactuals
@@
Questions that help to consider other possibilities ensuring that the larger solution space is considered. Contextual Information (e.g., differential diagnosis or patient preferences (Q7)), Partial Dependence Plot, Counterfactuals,
@@
Questions that elicit taken for granted assumptions to ensure that the decision-maker is aware of their reasons and can direct their actions accordingly. General Questions, Confidence of decision-maker
@@
Questions that help the decision-maker enquire and take into account the preferences or needs of the people concerned (e.g., patient), in order to ensure a better or broader understanding of the problem at hand. General questions, Preference Elicitation
Q8* Consequences of Decision
Questions that help elicit anticipation and forward-looking responsibility to ensure that (unintended) consequences are considered and mitigated. Limitations of Model, FactSheets (Arnold et al., 2019), Model Cards (Mitchell et al., 2019)
Q9* Change Intervention
Questions that support the exploration of feasible (smaller) interventions to make a desired outcome more likely. Feature Perturbation, Counterfactuals (Karimi et al., 2021; Poyiadzi et al., 2020)